# SynCo: Synthetic Hard Negatives in Contrastive Learning for Better Unsupervised Visual Representations

### v1

@@ Abstract
Contrastive learning has become a dominant approach in self-supervised visual representation learning, with hard negatives—samples that closely resemble the anchor—being key to enhancing the discriminative power of learned representations. However, efficiently leveraging hard negatives remains a challenge due to the difficulty in identifying and incorporating them without significantly increasing computational costs. To address this, we introduce SynCo (Synthetic negatives in Contrastive learning), a novel contrastive learning approach that improves model performance by generating synthetic hard negatives. Built on the MoCo framework, SynCo introduces six novel strategies for creating diverse synthetic hard negatives that can be generated on-the-fly with minimal computational overhead. SynCo achieves faster training and better representation learning, achieving a top-1 accuracy of 68.1% in ImageNet linear evaluation after only 200 epochs on pretraining, surpassing MoCo’s 67.5% with the same ResNet-50 encoder. Additionally, it transfers more effectively to detection tasks: on the PASCAL VOC, it outperforms both the supervised baseline and MoCo, achieving an $AP$ of 82.5%; on the COCO dataset, it sets a new benchmark with 40.4% $AP$ for bounding box detection and 35.4% $AP$ for instance segmentation. Our synthetic hard negative generation procedure significantly enhances the quality of visual representations learned through self-supervised contrastive learning. Code is available at https://github.com/giakoumoglou/synco.
@@ 1 Introduction
Figure 1: SynCo extends MoCo by introducing synthetic hard negatives generated on-the-fly from a memory queue. The process begins with two augmented views of an image, $\mathbf{x}_{q}$ and $\mathbf{x}_{k}$, processed by an encoder and a momentum encoder, respectively, producing feature vectors $\mathbf{q}$ and $\mathbf{k}$. The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{h}_{1},\mathbf{h}_{2},\ldots$ generated using the SynCo strategies. These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation. This approach enhances the learning process by providing more challenging and informative negative samples.
@@
Despite its efficacy, instance discrimination faces significant challenges. A key limitation is its dependence on numerous negative samples for optimal performance, often leading to increased computational costs. For example, SimCLR requires large batch sizes to provide sufficient negatives [6]. While approaches like MoCo mitigate some issues through a dynamic queue and momentum encoder [17, 8], they still confront the challenge of selecting and maintaining high-quality hard negatives—samples particularly similar to positives and thus more difficult to distinguish.
Recent studies [6, 10, 33, 39, 8, 20, 29, 3] highlight the importance of carefully crafted data augmentations in learning robust representations. These transformations likely provide more diverse, challenging copies of images, increasing the difficulty of the self-supervised task. This self-supervised task is a pretext problem (e.g., predicting image rotations [12] or solving jigsaw puzzles [27]) designed to induce learning of generalizable features without explicit labels. At the same time, data mixing techniques operating at either the pixel [46, 43] or the feature level [37] help models learn more robust features that improve both supervised and semi-supervised learning on subsequent (target) tasks.
The concept of hard negatives has been explored extensively as a means to enhance the robustness and generalization of contrastive learning models. Hard negatives are those that lie close to the decision boundary, making them crucial for fine-tuning the model’s discriminative capabilities. [20] introduced MoCHI, which integrates hard negative mixing into the MoCo framework, demonstrating significant improvements in performance. The task of efficiently creating and choosing hard negatives is important in contrastive self-supervised learning.
In this paper, we present SynCo (Synthetic negatives in Contrastive learning), a novel approach to contrastive learning that leverages synthetic hard negatives to enhance the learning process. Building on the foundations of MoCo, SynCo introduces six distinct strategies for generating synthetic hard negatives, each designed to provide diverse and challenging contrasts to the model. These strategies include: (1) Interpolated negatives; (2) Extrapolated negatives; (3) Mixup negatives; (4) Noise-injected negatives; (5) Perturbed negatives; and (6) Adversarial negatives. By incorporating these synthetic samples, SynCo aims to push the boundaries of contrastive learning, improving both the efficiency and effectiveness of the training process.
@@ 2.1 Contrastive Learning
Present contrastive learning methods learn representations by instance discrimination as a pretext task, treating each image as its own class. The core idea is to pull together an anchor and a ”positive” sample in the embedding space while pushing apart the anchor from ”negative” samples [6, 21]. Positive pairs are typically generated by creating multiple views of each data point [33], such as through luminance and chrominance decomposition [32], random augmentation [6, 17], image patches [36], or student-teacher model representations [13, 5, 28]. This approach achieves state-of-the-art performance on downstream tasks. The common training objective, provided by InfoNCE [36] or its variants [6, 10, 35], maximizes mutual information [18, 2], requiring numerous negative pairs. While large batch sizes [6] address this, they’re resource-intensive. Alternatively, MoCo [17, 8], PIRL [25], and InstDis [41] use memory structures. Recent advancements explore strategies like invariance regularizers [26], dataset-derived positives [10], and unified contrastive formulas [31]. Some approaches eliminate negative samples via asymmetric Siamese structures or normalization [13, 7, 5], while others prevent model collapse through redundancy reduction [44] or regularization [4]. Despite these innovations, instance discrimination’s fine-grained nature can lead to false-negative pairs [47].
@@ 2.2 Hard Negatives
In contrastive learning, using hard negatives is crucial for improving the learning process and model performance. Hard negatives are challenging yet relevant samples that help in defining the embedding space more effectively, as they allow the model to better distinguish between similar features. The use of hard negatives involves selecting negatives that are similar to positive samples but are different enough to aid the model in learning distinctive features. Dynamically sampling these hard negatives during training prevents the model from easily minimizing the loss, which enhances its learning capabilities [17, 6]. MoCo [17] uses a dynamic queue and a momentum-based encoder update mechanism. This method keeps the negatives fresh by periodically updating the queue with new samples, which maintains the challenge for the model throughout the training period. SimCLR [6] and InfoMin [33] suggest that the difficulty of negative samples can be adjusted by varying the data augmentation techniques used during training. This makes the task progressively harder, which is beneficial for training. Building on these ideas, MoCHI [20] extends the MoCo framework by introducing hard negative mixing into the embedding space.
When integrating these approaches into embedding space generation, it is essential to consider how negative samples impact the model’s learning space. The goal is to implement a strategy that not only provides a challenge to the model but also enhances its ability to extract and differentiate features effectively. By using these methods models become more capable of handling detailed and complex tasks. This approach ensures that each negative sample significantly contributes to the training, optimizing learning outcomes and boosting overall model effectiveness.
3 Hard Negatives in Contrastive Learning
3.1 Contrastive Learning with Memory
@@
Contrastive learning seeks to differentiate between similar and dissimilar data pairs, often treated as a dictionary look-up where representations are optimized to align positively paired data through contrastive loss in the embedding space [17]. Given an image $x$, and a distribution of image augmentation $\mathcal{T}$, we create two augmented views of the same image using the transformation $t_{q},t_{k}\sim\mathcal{T}$, i.e., $x_{q}=t_{q}(x)$ and $x_{k}=t_{k}(x)$. Two encoders, $f_{q}$ and $f_{k}$, namely the query and key encoders, generate the vectors $\mathbf{q}=f_{q}(x_{q})$ and $\mathbf{k}=f_{k}(x_{k})$ respectively. The learning objective minimizes a contrastive loss using the InfoNCE criterion [36]:
$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})=-\log\frac{\exp(\mathbf{q}\cdot\mathbf{k}/\tau)}{\exp(\mathbf{q}\cdot\mathbf{k}/\tau)+\sum\limits_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}\cdot\mathbf{n}/\tau)}$ (1)
@@
where $\mathbf{k}$ is $f_{k}$’s output from the same augmented image as $\mathbf{q}$, and $\mathcal{Q}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ includes outputs from different images, representing negative samples of size $K$. The temperature parameter $\tau$ adjusts scaling for the $\ell_{2}$-normalized vectors $\mathbf{q}$ and $\mathbf{k}$. The memory bank $\mathcal{Q}$ can be defined as an external memory containing every other image in the dataset [25, 32, 41], a queue of the last batches [17], or every other image in the current minibatch [6].
@@
The log-likelihood function of the contrastive loss in Equation 1 is defined over the probability distribution created by applying a softmax function for each input/query $\mathbf{q}$. The gradient of the loss with respect to the query $\mathbf{q}$ is given by:
@@
$p_{z_{i}}=\frac{\exp(\mathbf{q}\cdot\mathbf{z_{i}}/\tau)}{\sum_{j\in Z}\exp(\mathbf{q}\cdot\mathbf{z_{j}}/\tau)}$ (3)
with $\mathbf{z_{i}}$ being a member of the set $\mathcal{Q}\cup\{\mathbf{k}\}$. We see that the contributions of the positive and negative logits to the loss are identical to the ones for a $(K+1)$-way cross-entropy classification loss, where the logit for the key corresponds to the query’s latent class [1] and all gradients are scaled by $1/\tau$.
@@ 3.2 Understanding Hard Negatives
Hard negatives are critical for contrastive learning [1, 14, 19, 24, 41, 20]. Sampling negatives from the same batch leads to a need for larger batches [6] while sampling negatives from a memory bank that contains every other image in the dataset requires the time consuming task of keeping a large memory up-to-date [25, 41]. In the latter case, a trade-off exists between the ”freshness” of the memory bank representations and the computational overhead for re-computing them as the encoder keeps changing. MoCo [17] offers a compromise between the two negative sampling extremes: it keeps a queue of the latest $K$ features from the last batches, encoded with a second key encoder that trails the (main/query) encoder with a much higher momentum. For MoCo, the key feature $\mathbf{k}$ and all features in $\mathcal{Q}$ are encoded with the key encoder.
Hardness of negatives. Figure 2 illustrates the evolving difficulty of negatives during MoCo-v2 training on ImageNet-100. We plot the top 1024 matching probabilities $p_{z_{i}}$. Initially, at epoch 0, the logits exhibit a relatively uniform distribution. However, as training progresses, an increasing number of negatives cease to contribute significantly to the loss. This suggests that the majority of memory negatives become less effective for the proxy task over time. Previous works [17, 6] have shown that larger memory or batch sizes are necessary for obtaining more challenging negatives. Our analysis confirms that many memory negatives provide diminishing benefits to the learning process as training advances.
Figure 2: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where we rank the key over all negatives).
Difficulty of the proxy task. MoCo and SimCLR learn augmentation-invariant representations. We follow the hypothesis of [20] that proxy task difficulty correlates with transformation set complexity. Figure 3 shows the proxy task performance (percentage of queries ranking the key over all negatives) for MoCo and MoCo-v2 on ImageNet-100. MoCo-v2, despite lower proxy task performance due to harder augmentations, achieves better linear classification results compared to MoCo. This suggests that increasing proxy task difficulty through more complex augmentations can enhance overall performance.
Figure 3: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where we rank the key over all negatives).
@@ 4 Synthetic Hard Negatives in Contrastive Learning
Based on these observations, we present an approach for generating synthetic hard negatives in the embedding space using six distinct strategies. We refer to the proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
@@ 4.1 Hard Negatives Generation
Let $\mathbf{q}$ be the query image, $\mathbf{k}$ its key, and $\mathbf{n}\in\mathcal{Q}$ represent the negative features from a memory structure of size $K$. The loss for the query is calculated using logits $\ell(z_{i})=\mathbf{q}\cdot z_{i}/\tau$ fed into a softmax function. Define $\mathcal{\hat{Q}}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ as the ordered set of all negative features, such that $\ell(\mathbf{n}_{i})>\ell(\mathbf{n}_{j})$ for all $i<j$, meaning the set of negative features is sorted by decreasing similarity to the query feature. The hardest negatives are identified by truncating the ordered set $\mathcal{\hat{Q}}$, keeping only the first $N<K$ items, denoted as $\mathcal{\hat{Q}}^{N}$.
Interpolated hard negatives (type 1). For each query $\mathbf{q}$, we propose to generate $N_{1}$ hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{1}=\{\mathbf{h}_{1}^{1},\mathbf{h}_{2}^{1},\ldots,\mathbf{h}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
@@
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\alpha_{k}$ is randomly sampled from a uniform distribution in the range $(0,\alpha_{\text{max}})$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Interpolation creates a synthetic embedding that lies between the query and the negative in the embedding space. We set $\alpha_{\text{max}}=0.5$ to ensure that the query’s contribution is always smaller than the one of the negative. This is similar to the hardest negatives (type 2) of MoCHI.
Extrapolated hard negatives (type 2). For each query $\mathbf{q}$, we propose to generate $N_{2}$ hard negative features by extrapolating beyond the query embedding in the direction of the hardest negative features. Similar to the interpolated method, we use a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{2}=\{\mathbf{h}_{1}^{2},\mathbf{h}_{2}^{2},\ldots,\mathbf{h}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
@@
Mixup hard negatives (type 3). For each query $\mathbf{q}$, we propose to generate $N_{3}$ hard negative features by combining pairs of the $N$ hardest existing negative features in $\mathcal{\hat{Q}}^{N}$. Let $H^{3}=\{\mathbf{h}_{1}^{3},\mathbf{h}_{2}^{3},\ldots,\mathbf{h}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
@@
where $\mathbf{n}_{i},\mathbf{n}_{j}\in\mathcal{\hat{Q}}^{N}$ and $\gamma_{k}$ is randomly sampled from a uniform distribution in the range $(0,1)$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Mixup combines pairs of the hardest existing negative features to create a synthetic embedding that represents a blend of challenging cases. This is similar to the hard negatives (type 1) of MoCHI.
Noise-injected hard negatives (type 4). For each query $\mathbf{q}$, we propose to generate $N_{4}$ hard negative features by adding Gaussian noise to the hardest negative features. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{4}=\{\mathbf{h}_{1}^{4},\mathbf{h}_{2}^{4},\ldots,\mathbf{h}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(0,\sigma^{2})$ (7)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise with standard deviation $\sigma$. The noisy negatives are normalized before being used in the loss calculation. Noise injection adds Gaussian noise to the hardest negative features, resulting in a synthetic embedding with added randomness.
Perturbed hard negatives (type 5). For each query $\mathbf{q}$, we propose to generate $N_{5}$ hard negative features by perturbing the embeddings of the hardest negative features. Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{5}=\{\mathbf{h}_{1}^{5},\mathbf{h}_{2}^{5},\ldots,\mathbf{h}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by:

### v3

@@ Abstract
Contrastive learning has become a dominant approach in self-supervised visual representation learning. Hard negatives - samples closely resembling the anchor - are key to enhancing learned representations’ discriminative power. However, efficiently leveraging hard negatives remains challenging. We introduce SynCo (Synthetic negatives in Contrastive learning), a novel approach that improves model performance by generating synthetic hard negatives on the representation space. Building on the MoCo framework, SynCo introduces six strategies for creating diverse synthetic hard negatives on-the-fly with minimal computational overhead. SynCo achieves faster training and better representation learning, reaching 67.9% top-1 accuracy on ImageNet ILSVRC-201 linear evaluation after 200 pretraining epochs, surpassing MoCo’s 67.5% using the same ResNet-50 encoder. It also transfers more effectively to detection tasks: on PASCAL VOC, it outperforms both the supervised baseline and MoCo with 82.6% AP; on COCO, it sets new benchmarks with 41.0% AP for bounding box detection and 35.7% AP for instance segmentation. Our synthetic hard negative generation approach significantly enhances visual representations learned through self-supervised contrastive learning. Code is available at https://github.com/giakoumoglou/synco.
@@ 1 Introduction
Figure 1: SynCo extends MoCo by introducing synthetic hard negatives generated on-the-fly from a memory queue. The process begins with two augmented views of an image, $\mathbf{x}_{q}$ and $\mathbf{x}_{k}$, processed by an encoder and a momentum encoder, respectively, producing feature vectors $\mathbf{q}$ and $\mathbf{k}$. The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{h}_{1},\mathbf{h}_{2},\ldots$ generated using the SynCo strategies. These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation.
@@
Despite its effectiveness, instance discrimination faces challenges. A key limitation is the need for numerous negative samples, often leading to increased computational costs. For example, SimCLR requires large batch sizes for sufficient negatives [6]. While approaches like MoCo address some issues through dynamic queues and momentum encoders [20, 8], they still face challenges in selecting and maintaining high-quality hard negatives.
Recent studies have highlighted the importance of carefully crafted data augmentations in learning robust representations [6, 11, 39, 45, 8, 23, 34, 3]. These transformations likely provide more diverse, challenging copies of images, increasing the difficulty of the self-supervised task. This self-supervised task is a pretext problem (e.g., predicting image rotations [13] or solving jigsaw puzzles [32]) designed to induce learning of generalizable features without explicit labels. Moreover, techniques that combine data at the pixel level [53, 50] or at the feature level [43] have proven effective in helping models learn more resilient features, leading to improvements in both fully supervised and semi-supervised tasks.
The concept of challenging negative samples has been explored as a way to enhance contrastive learning models. These samples, which lie close to the decision boundary, are crucial for refining the model’s discriminative abilities. Recent work like MoCHI [23] has shown improvements by incorporating harder negatives. Our work builds on this idea by proposing novel strategies for synthetic hard negative generation.
In this paper, we present SynCo (Synthetic negatives in Contrastive learning), a novel approach to contrastive learning that leverages synthetic hard negatives to enhance the learning process. Building on the foundations of MoCo, SynCo introduces six distinct strategies for generating synthetic hard negatives, each designed to provide diverse and challenging contrasts to the model. These strategies include: interpolated negatives; extrapolated negatives; mixup negatives; noise-injected negatives; perturbed negatives; and adversarial negatives. By incorporating these synthetic samples, SynCo aims to push the boundaries of contrastive learning, improving both the efficiency and effectiveness of the training process.
@@ 2.1 Contrastive Learning
Recent contrastive learning methods focus on instance discrimination as a pretext task, treating each image as its own class. The core principle involves bringing an anchor and a ”positive” sample closer in the representation space while pushing the anchor away from ”negative” samples [6, 24]. Positive pairs are typically created through multiple views of each data point [39], using techniques such as color decomposition [38], random augmentation [6, 20], image patches [42], or student-teacher model representations [14, 5, 33]. The common training objective, based on InfoNCE [42] or its variants [6, 11, 41], aims to maximize mutual information [21, 2], necessitating numerous negative pairs. While some approaches use large batch sizes [6] to address this, others like MoCo [20, 8], PIRL [30], and InstDis [47] employ memory structures. Recent advancements explore strategies such as invariance regularizers [31], dataset-derived positives [11], and unified contrastive formulas [37]. Some methods eliminate negative samples through asymmetric Siamese structures or normalization [14, 7, 5], while others prevent model collapse via redundancy reduction [51] or regularization [4]. Despite these innovations, the fine-grained nature of instance discrimination can lead to false-negative pairs [55].
@@ 2.2 Hard Negatives
Hard negatives are critical in contrastive learning as they improve the quality of visual representations by helping to define the representation space more effectively. These challenging yet relevant samples are harder to distinguish from the anchor point, enabling the model to better differentiate between similar features. The use of hard negatives involves selecting samples that are similar to positive samples but different enough to aid in learning distinctive features. Dynamic sampling of hard negatives during training prevents the model from easily minimizing the loss, enhancing its learning capabilities [20, 6]. Various approaches have been proposed to leverage hard negatives effectively. For instance, MoCo [20] utilizes a dynamic queue and momentum-based encoder updates to maintain fresh and challenging negatives throughout training. Other methods, such as SimCLR [6] and InfoMin [39], suggest adjusting the difficulty of negative samples by varying data augmentation techniques. This progressive increase in task difficulty benefits the training process. Building on these ideas, MoCHI [23] has explored integrating hard negative mixing into existing frameworks to further improve performance. When incorporating these approaches into representation space generation, it’s essential to consider how negative samples impact the model’s learning space. The goal is to implement a strategy that not only challenges the model but also enhances its ability to extract and differentiate features effectively. By employing these methods, models become more adept at handling detailed and complex tasks, ensuring each negative sample significantly contributes to optimizing learning outcomes and boosting overall model effectiveness.
2.3 Synthetic Features
Synthetic feature generation is a widely used method to enhance deep learning models, especially in cases with limited labeled data. By adding synthetic features to the representation space, models improve in generalization and performance. Methods have generated features for unseen classes using generative models [17, 48, 36], while others have integrated these into self-supervised and contrastive learning frameworks [27, 54]. Models that combine generation with representation spaces have also shown success in zero-shot learning [16]. In contrast, our approach directly generates synthetic hard negatives in contrastive learning, improving representation without additional generative models.
3 Preliminaries
3.1 Contrastive Learning
@@
Contrastive learning seeks to differentiate between similar and dissimilar data pairs, often treated as a dictionary look-up where representations are optimized to align positively paired data through contrastive loss in the representation space [20]. Given an image $x$, and a distribution of image augmentation $\mathcal{T}$, we create two augmented views of the same image using the transformation $t_{q},t_{k}\sim\mathcal{T}$, i.e., $x_{q}=t_{q}(x)$ and $x_{k}=t_{k}(x)$. Two encoders, $f_{q}$ and $f_{k}$, namely the query and key encoders, generate the vectors $\mathbf{q}=f_{q}(x_{q})$ and $\mathbf{k}=f_{k}(x_{k})$ respectively. The learning objective minimizes a contrastive loss using the InfoNCE criterion [42]:
$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})=-\log\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)}{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)+\sum\limits_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}^{T}\cdot\mathbf{n}/\tau)}$ (1)
@@
Here, $\mathbf{k}$ is $f_{k}$’s output from the same augmented image as $\mathbf{q}$, and $\mathcal{Q}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ includes outputs from different images, representing negative samples of size $K$. The temperature parameter $\tau$ adjusts scaling for the $\ell_{2}$-normalized vectors $\mathbf{q}$ and $\mathbf{k}$. The key encoder $f_{k}$ can be updated in two ways. In the synchronized update approach, $f_{k}$ is updated synchronously with $f_{q}$, maintaining identical weights throughout training [6]. Alternatively, a momentum update scheme can be employed, where $f_{k}$ is updated using the equation: $\theta_{k}\leftarrow m\cdot\theta_{k}+(1-m)\cdot\theta_{q}$ [20]. Here, $\theta_{k}$ and $\theta_{q}$ are the parameters of $f_{k}$ and $f_{q}$ respectively, and $m\in[0,1]$ is the momentum coefficient. This momentum approach allows $f_{k}$ to evolve more slowly, providing more consistent negative samples over time and potentially stabilizing the learning process. The memory bank $\mathcal{Q}$ can be defined in various ways, such as an external memory of all dataset images [30, 38, 47], a queue of recent batches [20], or the current minibatch [6].
@@
The gradient of the contrastive loss in Equation 1 with respect to the query $\mathbf{q}$ is given by:
@@
$p_{z_{i}}=\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{z_{i}}/\tau)}{\sum_{j\in Z}\exp(\mathbf{q}^{T}\cdot\mathbf{z_{j}}/\tau)}$ (3)
with $\mathbf{z_{i}}$ being a member of the set $\mathcal{Q}\cup\{\mathbf{k}\}$. The positive and negative logits contribute to the loss similarly to a $(K+1)$-way cross-entropy classification, with the key logit representing the query’s latent class [1].
@@ 3.2 Understanding Hard Negatives
The effectiveness of contrastive learning approaches hinges critically on the utilization of hard negatives [1, 15, 22, 29, 47, 23]. Current approaches face significant challenges in efficiently leveraging these hard negatives. Sampling from within the same batch necessitates larger batch sizes [6, 9], potentially straining computational resources. Conversely, maintaining a memory bank containing representations of the entire dataset incurs substantial computational overhead in keeping the memory up-to-date [30, 47, 20, 8]. These limitations underscore the need for more efficient strategies to generate and utilize hard negatives in contrastive learning frameworks.
Hardness of negatives. The ”hardness” of negative samples, defined by their similarity to positive samples in the representation space, determines how challenging they are for the model to differentiate, directly impacting the effectiveness of the contrastive learning process. Figure 2 illustrates the evolution of negative sample hardness during MoCo-v2 training on ImageNet-100. The plot depicts the top 1024 matching probabilities $p_{z_{i}}$ across different training epochs. Initially, the distribution of these probabilities is relatively uniform. However, as training progresses, a clear trend emerges: fewer negatives contribute significantly to the loss function. This observation suggests that the model rapidly learns to distinguish most negatives, leaving only a small subset that remains challenging. Such a phenomenon underscores the importance of maintaining a diverse pool of hard negatives throughout the training process to sustain effective learning [23].
Figure 2: Histogram of the top 1024 matching probabilities $p_{z_{i}}$, $z_{i}\in\mathcal{Q}$ for MoCo-v2, over various epochs. Logits are organized in descending order, and each line indicates the mean matching probability across all queries [23].
Difficulty of the proxy task. The difficulty of the proxy task in contrastive learning, typically defined by the self-supervised objective, significantly influences the quality of learned representations. Figure 3 compares the proxy task performance of MoCo and MoCo-v2 on ImageNet-100, measured by the percentage of queries where the key ranks above all negatives. Notably, MoCo-v2, which employs more aggressive augmentations, exhibits lower proxy task performance compared to MoCo, indicating a more challenging learning objective. Paradoxically, this increased difficulty correlates with improved performance on downstream tasks such as linear classification [23]. This counterintuitive relationship between proxy task difficulty and downstream performance suggests that more challenging self-supervised objectives can lead to the learning of more robust and transferable representations, motivating the development of strategies to dynamically modulate task difficulty during training.
Figure 3: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where the key is ranked higher than all negatives).
@@ 4 Synthetic Hard Negatives in Contrastive Learning
Based on these observations, we present an approach for generating synthetic hard negatives in the representation space using six distinct strategies. We refer to the proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
@@ 4.1 Hard Negatives Generation
Let $\mathbf{q}$ represent the query image, $\mathbf{k}$ its corresponding key, and $\mathbf{n}\in\mathcal{Q}$ denote the negative features from a memory structure of size $K$. The loss associated with the query is computed using the logits $\ell(\mathbf{z_{i}})=\mathbf{q}^{T}\cdot\mathbf{z_{i}}/\tau$, which are processed through a softmax function. We define $\mathcal{\hat{Q}}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ as the ordered set of all negative features, where $\ell(\mathbf{n}_{i})>\ell(\mathbf{n}_{j})$ for all $i<j$, implying that the negative features are sorted based on decreasing similarity to the query. The most challenging negatives are selected by truncating the ordered set $\mathcal{\hat{Q}}$, retaining only the first $N<K$ elements, denoted as $\mathcal{\hat{Q}}^{N}$.
Interpolated hard negatives (type 1). For each query $\mathbf{q}$, we propose to generate $N_{1}$ synthetic hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{1}=\{\mathbf{h}_{1}^{1},\mathbf{h}_{2}^{1},\ldots,\mathbf{h}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
@@
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\alpha_{k}$ is randomly sampled from a uniform distribution in the range $(0,\alpha_{\text{max}})$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Interpolation creates a synthetic embedding that lies between the query and the negative in the representation space. We set $\alpha_{\text{max}}=0.5$ to guarantee that the contribution of the query is always less than that of the negative. This is similar to the hardest negatives (type 2) of MoCHI [23].
Extrapolated hard negatives (type 2). For each query $\mathbf{q}$, we propose to generate $N_{2}$ hard negative features by extrapolating beyond the query embedding in the direction of the hardest negative features. Similar to the interpolated method, we use a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{2}=\{\mathbf{h}_{1}^{2},\mathbf{h}_{2}^{2},\ldots,\mathbf{h}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
@@
Mixup hard negatives (type 3). For each query $\mathbf{q}$, we propose to generate $N_{3}$ hard negative features by combining pairs of the $N$ hardest existing negative features in $\mathcal{\hat{Q}}^{N}$. Let $H^{3}=\{\mathbf{h}_{1}^{3},\mathbf{h}_{2}^{3},\ldots,\mathbf{h}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
@@
where $\mathbf{n}_{i},\mathbf{n}_{j}\in\mathcal{\hat{Q}}^{N}$ and $\gamma_{k}$ is randomly sampled from a uniform distribution in the range $(0,1)$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Mixup combines pairs of the hardest existing negative features to create a synthetic embedding that represents a blend of challenging cases. This is similar to the hard negatives (type 1) of MoCHI [23].
Noise-injected hard negatives (type 4). For each query $\mathbf{q}$, we propose to generate $N_{4}$ hard negative features by adding Gaussian noise to the hardest negative features. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{4}=\{\mathbf{h}_{1}^{4},\mathbf{h}_{2}^{4},\ldots,\mathbf{h}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ (7)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ represents Gaussian noise with standard deviation $\sigma$ (where $\mathbf{I}$ is the identity matrix). The noisy negatives are normalized before being used in the loss calculation. Noise injection adds Gaussian noise to the hardest negative features, resulting in a synthetic embedding with added randomness.
Perturbed hard negatives (type 5). For each query $\mathbf{q}$, we propose to generate $N_{5}$ hard negative features by perturbing the embeddings of the hardest negative features. Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{5}=\{\mathbf{h}_{1}^{5},\mathbf{h}_{2}^{5},\ldots,\mathbf{h}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by: